OpenStudy (ciarabelen):

Which sets of coordinates is a solution of the equation 2x-y=11? A.) (-6,-1) B.) (-1,9) C.) (0,11) D.) (2,-7) Can you explain the answer to me?

OpenStudy (whpalmer4):

Well, you can try them out and see if they produce a true statement. For example, let's try answer C: \[2(0)-11 = 11\]\[-11=11\]That's not true, so (0,11) is not a solution to that equation.

OpenStudy (whpalmer4):

How about (8,5)? \[2(8)-5=11\]\[16-5=11\]\[11=11\]That is true, so (8,5) would be a solution (though it does not appear on your answer list).

pooja195 (pooja195):

2x -y = 11 -y = -2x + 11 y = 2x -11 im thinking d but not sure

OpenStudy (ciarabelen):

It was D.) Thanks :)

pooja195 (pooja195):

np